A pleasant day! please help in this one.
find a quadratic equation with integral coefficients whose roots are the reciprocals of the roots of:
x^2 + 2x - 15 = 0
Thank you very much!!!

Answer by sachi(548)

x^2 + 2x - 15 = 0
or x^2 + 5x-3x - 15 = 0
or x[x+5]-3[x+5]=0
or [x+5][x-3]=0
or x=-5 or 3
new roots are -1/5 & 1/3
so the eqn is x^2-x[-1/5+1/3]+{-1/5}*1/3
or x^2-2x/15-1/15=0
or 15x^2-2x-1=0
ans